You are riding a unicorn at 25 m/s and come to a uniform stop at a red light 20 m away what s your acceleration?
1
Expert's answer
2021-10-13T15:40:56-0400

"v_f^2 -v_i^2 = 2ax \\
0 -25^2 = 2a \times 20 \\
a = \frac{-625}{40} = -15.625 \; m/s^2"